A 95% confidence interval for μ, the true mean city gas mileage for a particular vehicle is (23.0, 26.0). The sample mean, x¯, and margin of error for this interval are:
Answer #1

Solution:

Given , 95% confidence interval for mean is  (23.0, 26.0)

Upper limit = 26.0

Lower Limit = 23.0

Bur ,

Upper limit = Sample mean + Margin of error

Lower Limit = Sample mean - Margin of error

Solving we get ,

Sample mean = (Upper Limit + Lower Limit)/2 = (26.0 + 23.0)/2 = 24.5

Margin of error = (Upper Limit - Lower Limit)/2 = (26.0 - 23.0)/2 = 1.5

Answer :

Sample mean = 24.5

Margin of error = 1.5